A sample weighing 4.52g containing KHP (MM=204.2 g/mol) required 28 ml of 0.15M NaOH solution for neutralization, calculate the mass percent of KHP in this sample
1
Expert's answer
2020-05-25T14:13:21-0400

Solution:

KHP = potassium hydrogen phthalate = KHC8H4O4


The balanced chemical equation:

NaOH(aq) + KHP(aq) = NaKP(aq) + H2O(l)

According to the chemical equation: n(NaOH) = n(KHP)


Moles of NaOH = n(NaOH) = Molarity of NaOH × Volume of solution (NaOH)

n(NaOH) = (0.15 M × 0.028 L) = 0.0042 mol


n(KHP) = n(NaOH) = 0.0042 mol

Moles of KHP = n(KHP) = Mass of KHP / Molar mass of KHP

Mass of KHP = m(KHP) = Moles of KHP × Molar Mass of KHP

m(KHP) = (0.0042 mol × 204.2 g/mol) = 0.85764 g


% KHP = [m(KHP) / m(sample)] × 100%

% KHP = (0.85764g / 4.52 g) × 100% = 0.1897 × 100% = 18.97%

%KHP = 18.97 %


Answer: The mass percent of KHP in this sample is 18.97%